Question 14

XYZ Company uses the formula y = a + bX to predict and analyze overhead costs. In the previous year, XYZ used $1,750 per month for the a factor and $0.35 for the b factor in applying overhead. XYZ has used direct labour hours in the past, but is wondering whether overhead behaviour is more closely associated with machine hours. The following data have been generated for consideration:

Month Machine Hours Overhead Costs
1 414 $2,752
2 487 $2,968
3 394 $2,436
4 483 $2,824
5 510 $2,745
6 443 $2,693

a) Determine the fixed and variable values, using machine hours.

b)If it is expected that the company will incur 1,032 machine hours in the coming month, predict that month’s overhead costs.  

c)or April of the coming year, XYZ expects 3,300 direct labour hours and 604 machine hours. Predict overhead costs using

direct labour hours and machine hours
Answer #1

Ansever - Page No ③ XYZ Company uses the formula ya a+bX In previous gear, fixed corf - $1450 per month Variable cook= $0.35Page No ③ 6 Months verhead coat Company in cur machine hours 1032 Months overhead cosf = 1388 + (1032 x $2.66) = 1388 + 2745